# Two-Stage Pretraining for Molecular Property Prediction in the Wild

### v1

@@ Abstract
Accurate property prediction is crucial for accelerating the discovery of new molecules. Although deep learning models have achieved remarkable success, their performance often relies on large amounts of labeled data that are expensive and time-consuming to obtain. Thus, there is a growing need for models that can perform well with limited experimentally-validated data. In this work, we introduce MoleVers, a versatile pretrained model designed for various types of molecular property prediction in the wild, i.e., where experimentally-validated molecular property labels are scarce. MoleVers adopts a two-stage pretraining strategy. In the first stage, the model learns molecular representations from large unlabeled datasets via masked atom prediction and dynamic denoising, a novel task enabled by a new branching encoder architecture. In the second stage, MoleVers is further pretrained using auxiliary labels obtained with inexpensive computational methods, enabling supervised learning without the need for costly experimental data. This two-stage framework allows MoleVers to learn representations that generalize effectively across various downstream datasets. We evaluate MoleVers on a new benchmark comprising 22 molecular datasets with diverse types of properties, the majority of which contain 50 or fewer training labels reflecting real-world conditions. MoleVers achieves state-of-the-art results on 20 out of the 22 datasets, and ranks second among the remaining two, highlighting its ability to bridge the gap between data-hungry models and real-world conditions where practically-useful labels are scarce.
@@ 1 Introduction
A reliable molecular property prediction model enables researchers to efficiently screen vast numbers of potential compounds, reducing the need for costly experimental validation to only a select few. To this end, deep learning-based approaches have demonstrated remarkable accuracy in predicting a range of molecular properties including electronic, physical, and bioactivity properties (Yang et al., 2019; Rong et al., 2020; Fang et al., 2022; Zhou et al., 2023). However, these models typically rely on large datasets with hundreds of thousands of labeled data to achieve strong predictive performance. Yet, in real-world scenarios, labeled molecular data is often limited. For example, of the 1,644,390 assays in the ChemBL database (Zdrazil et al., 2024), only 6,113 (0.37%) contain 100 or more labeled molecules. This raises doubts about whether existing models are suitable for molecular property prediction in the wild, i.e., in real-world scenarios where experimentally-validated data is scarce.
In this work, we introduce MoleVers, a versatile pretrained model designed for molecular property prediction in data-scarce scenarios. MoleVers is pretrained in two stages to maximize its generalizability to various types of downstream properties. In the first stage of pretraining, we employ masked atom prediction (MAP) and dynamic denoising with relatively large noise scales to improve the generalizability of the learned representations. While previous studies have shown that increasing noise scales can impair training stability and downstream performance (Zhou et al., 2023; Yang et al., 2024), MoleVers overcomes this issue through a novel branching encoder architecture that decouples the MAP and denoising pipelines. In the second pretraining stage, MoleVers refines its representations by predicting auxiliary properties that can be derived from inexpensive computational methods such as the Density Functional Theory (DFT). Given that molecular properties are often related to molecular structures, representations learned for predicting one property can also be useful for others. As a result of the two-stage pretraining, the model can learn molecular representations that improve the performance in downstream datasets.
@@
To evaluate MoleVers, we introduce a new benchmark, Molecular Property Prediction in the Wild (MPPW), that consists of 22 small datasets curated from the ChemBL database (Zdrazil et al., 2024). These datasets, most of which contain 50 or fewer training labels, span a wide range of molecular properties from physical characteristics to biological activities. We standardized the pretraining datasets and data splits to ensure fair comparisons between MoleVers and several state-of-the-art pretrained models. Experimental results show that MoleVers outperforms all baselines in 20 out of the 22 assays and ranks a close second in the remaining two, while no baseline consistently ranks in the top two. Moreover, MoleVers achieves state-of-the-art performance on large datasets in the MoleculeNet benchmark (Wu et al., 2018), highlighting the effectiveness of our two-stage pretraining strategy.
@@
In summary, our contributions are: (1) a two-stage pretraining framework that includes a novel dynamic denoising pretraining for learning molecular representations without requiring additional downstream labels, (2) a branching encoder that facilitates denoising pretraining with larger noise scales, and (3) the MPPW benchmark, designed to reflect real-world data limitations. All relevant source code will be publicly available upon publication.
2 Related Works
@@
Deep learning-based molecular property prediction has demonstrated remarkable successes. Early approaches use graph neural networks (GNNs) to learn molecular representations directly from molecular structures (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018). GNNs typically learn molecular representations by updating the node (atom) and edge (bond) features through a series of message passing accross neighboring atoms. Recently, popular property prediction benchmarks such as MoleculeNet (Wu et al., 2018) are dominated by transformer-based models (Luo et al., 2022; Zhou et al., 2023; Yang et al., 2024) that leverage self-attention mechanisms to learn long-range interactions between atoms in a molecule.
@@
Parallel to architectural advancements, pretraining has emerged as an effective strategy to improve property prediction peformance when labeled data is limited. By pretraining on a large, unlabeled dataset, a model can learn robust and transferable molecular representations that generalize well to a variety of downstream tasks. Various pretraining strategies have been proposed, including masked predictions (Wang et al., 2019; Xia et al., 2023; Zhou et al., 2023; Yang et al., 2024) and contrastive learning (Liu et al., 2022; Xia et al., 2023; Wang et al., 2022). Additionally, denoising atom coordinates and pairwise distance (Zaidi et al., 2023; Zhou et al., 2023; Liu et al., 2023) have been shown to lead to strong downstream performance. Denoising pretraining is equivalent to learning an approximate molecular force field (Zaidi et al., 2023; Liu et al., 2023), which could explain its effectiveness for improving downstream property prediction performance.
@@ 3 Two-Stage Pretraining
Our primary objective is to obtain an accurate molecular property prediction model without the need for additional, difficult-to-acquire labels for downstream tasks. To address this challenge, we propose a two-stage pretraining framework specifically designed to improve the generalization capability of our model, MoleVers. This approach enables accurate property prediction while minimizing the need for downstream labels during finetuning.
3.1 Stage 1: Masked Atom Prediction and Dynamic Denoising
The properties of a molecule are strongly influenced by the spatial arrangement of its atoms in the three-dimensional (3D) space. Consequently, self-supervised pretraining that involves both atom types and 3D structures is crucial for achieving strong performance in the downstream datasets. In the first stage of our pretraining framework, we employ masked atom prediction (MAP) and dynamic denoising to train MoleVers on a large, unlabeled dataset. This encourages the model to learn representations that are transferable to downstream datasets.
@@ 3.1.1 Masked Atom Prediction
Inspired by masked token prediction in natural language processing (NLP) (Devlin et al., 2019; Liu et al., 2019; Lewis et al., 2020), masked atom prediction (MAP) involves training a model to predict the correct atom types in a partially-masked molecule. This encourages the model to learn contextual relationship between atom types, capturing how they co-exist in various molecules. Multiple works (Zhou et al., 2023; Xia et al., 2023; Yang et al., 2024) have demonstrated the effectiveness of MAP as a pretraining task, which ultimately leads to better prediction models for the downstream datasets.
3.1.2 Dynamic Denoising
@@
We hypothesize that using a dynamic noise scale with larger values, e.g., drawn from a uniform distribution $\sigma\sim\mathcal{U}(a,b)$, where $b>1$, could improve the generalization ability of the model. Increasing $\sigma$ broadens each Gaussian distribution, allowing the learned force field to better cover molecules not seen in ${\mathbb{D}}^{\text{train}}$. Alternatively, dynamic denoising can be viewed as an augmentation technique. This follows the simple intuition that a larger $\sigma$ exposes the model to a wider set of non-equilibrium molecular configuration in a similar way to how diffusion models are trained (Ho et al., 2020; Song et al., 2021).
3.1.3 Decoupling MAP from Denoising
@@
Unfortunately, previous works have found that setting $\sigma$ to larger values often reduces the pretraining quality and downstream performance (Zhou et al., 2023; Yang et al., 2024). This phenomenon can be explained by comparing the complexities of the MAP and denoising tasks during pretraining. In MAP, the model learns to map masked atoms (${\bm{A}}^{\text{mask}}$) to their corresponding atom logits ($\hat{\bm{A}}$), $f({\bm{A}}^{\text{mask}})=\hat{\bm{A}}$, while in coordinate denoising, it learns to map noisy coordinates to their pristine values, $g(\tilde{{\bm{P}}})=\hat{\bm{P}}$. The MAP function is relatively simpler because it maps a finite set of inputs (atom types) to a relatively compact set of outputs (softmax-normalized logits). In contrast, denoising deals with continuous input and output coordinates, making it more complex as the number of possible mappings is much larger. When a single model handles both MAP and denoising, the overall complexity is dominated by the more challenging denoising task. The downstream performance could then be negatively affected if the model struggles to accurately fit the complex denoising function.
@@
This motivates us to introduce a branching encoder architecture, shown in Figure 1, that decouples the MAP and denoising pipelines. The branching design ensures that the complexity of the MAP task is minimally affected by the denoising. Furthermore, we propose to connect the two encoders with an aggregator module so that information can flow between the two pipelines.
@@
Figure 2: In pretraining stage 2 and finetuning, we keep only the primary encoder to encode the atom and pair distance representations. A prediction head is appended to the model to predict the properties of the molecules.
3.1.4 Branching Encoder
@@
Inspired by prior works in NLP that have found masked prediction to often be the most effective pretraining tasks (Lewis et al., 2020; Raffel et al., 2020), we set the MAP encoder as the primary encoder of the model. The primary encoder, shown in Figure 1, will be passed to the second pretraining stage (Figure 2 and used for predictions in the downstream datasets.
@@
The branching encoder takes as inputs the types ${\bm{A}}\in{\mathbb{Z}}^{N}$ and coordinates ${\bm{P}}\in{\mathbb{R}}^{N\times 3}$ of the $N$ atoms in a molecule. Following Zhou et al. (2023) and Yang et al. (2024), each atom type is encoded into atom representation ${\bm{X}}\in{\mathbb{R}}^{N\times C}$ where $C$ is the number of features, and the coordinates are transformed into pair distance representation ${\bm{D}}\in{\mathbb{R}}^{N\times N}$. During the first pretraining stage, the atom representations are masked with a ratio of r. We denote the masked atom representations as ${\bm{X}}^{m}$.
@@
To extract the molecule representation ${\bm{F}}$, we feed ${\bm{X}}^{m}$ and ${\bm{D}}$ into the primary MAP encoder $\phi^{p}$ The logits $\hat{\bm{A}}$ that represent the pristine atom types are then predicted with the MAP head $\psi^{p}$,
@@
$\tilde{{\bm{X}}}={\bm{X}}+\bm{\epsilon}_{1},\hskip 10.00002pt\tilde{{\bm{P}}}={\bm{P}}+\bm{\epsilon}_{2},\hskip 10.00002pt\bm{\epsilon}_{1}\sim\mathcal{N}(\textbf{0},\sigma^{2}{\bm{I}}_{N\times 1}),\hskip 10.00002pt\bm{\epsilon}_{2}\sim\mathcal{N}(\textbf{0},\sigma^{2}{\bm{I}}_{N\times 3}),\hskip 10.00002pt\sigma\sim\mathcal{U}(1,3).$ (2)
@@
$\tilde{{\bm{X}}}^{\text{aug}}=\text{concat}(\tilde{{\bm{X}}},\phi^{a}({\bm{F}})),\hskip 15.00002pt{\bm{G}}=\phi^{d}(\tilde{{\bm{X}}}^{\text{aug}},\tilde{{\bm{D}}}),\hskip 15.00002pt(\hat{\bm{X}},\hat{\bm{P}},\hat{\bm{D}})=\psi^{d}({\bm{G}}),$ (3)

### v2

@@ Abstract
Molecular deep learning models have achieved remarkable success in property prediction, but they often require large amounts of labeled data. The challenge is that, in real-world applications, labels are extremely scarce, as obtaining them through laboratory experimentation is both expensive and time-consuming. In this work, we introduce MoleVers, a versatile pretrained molecular model designed for various types of molecular property prediction in the wild, i.e., where experimentally-validated labels are scarce. MoleVers employs a two-stage pretraining strategy. In the first stage, it learns molecular representations from unlabeled data through masked atom prediction and extreme denoising, a novel task enabled by our newly introduced branching encoder architecture and dynamic noise scale sampling. In the second stage, the model refines these representations through predictions of auxiliary properties derived from computational methods, such as the density functional theory or large language models. Evaluation on 22 small, experimentally-validated datasets demonstrates that MoleVers achieves state-of-the-art performance, highlighting the effectiveness of its two-stage framework in producing generalizable molecular representations for diverse downstream properties.
Machine Learning, ICML
@@ 1 Introduction
Experimental chemistry is the domain of small data. Determining the property of a molecule requires carefully designed experiments, constrained by material costs, reaction time, and specialized infrastructure. This inherent scarcity forces chemists to extract insights from limited data points using statistical models, computational tools, and expert intuition, which in turn can be used to predict the properties of a novel molecule. With recent advances in deep learning, researchers have increasingly turned to deep neural networks for learning molecular property prediction directly from data (Yang et al., 2019; Rong et al., 2020; Fang et al., 2022; Zhou et al., 2023). However, these models typically rely on large datasets, where thousands or more labeled molecules are available for training. Contrasts this to real-world datasets; of the 1,644,390 assays in the ChemBL database (Zdrazil et al., 2024), only 6,113 assays (0.37%) contain 100 or more labeled molecules. This discrepancy poses a challenge: how can we adapt data-hungry deep learning models to real-world applications, where even 50 training labels are considered plenty?
One promising solution is through unsupervised pretraining on large, unlabeled molecular datasets, followed by finetuning on specific properties using few labeled molecules. Recent studies have explored this approach (Liu et al., 2022; Xia et al., 2023; Zhou et al., 2023; Yang et al., 2024), demonstrating improvements on standard molecular benchmarks. However, these benchmarks, such as MoleculeNet (Wu et al., 2018), contain thousands of labeled molecules—far from the reailty of most chemistry datasets. When evaluated on small datasets, these pretrained models often show a significant drop in performance, sometimes only marginally surpassing the baseline of simply predicting the mean values of the labels. Clearly, we need to rethink pretraining strategies in order to develop molecular property prediction models that are effective in real-world, small-data regime.
In this work, we introduce MoleVers, a versatile pretrained model designed for molecular property prediction in data-scarce scenarios. MoleVers is pretrained in two stages to maximize its generalizability to various types of downstream properties. In the first pretraining stage, we propose a joint masked atom prediction (MAP) and extreme denoising enabled by a novel branching encoder architecture and dynamic scale sampling. The extreme denoising approach allows the model to learn from a more diverse set of non-equilibrium configurations by utilizing a wider range of noise scales compared with conventional coordinate-based denoising. In the second pretraining stage, we train MoleVers to predict auxiliary properties that are different from, but more straightforward to obtain than the downstream properties. These auxiliary properties can be calculated using density functional theory (DFT) or large language models (LLMs). We use DFT to calculate three auxiliary properties: HOMO, LUMO, and dipole moment. For LLM-based auxiliary labels, we propose to use pairwise rankings of molecular properties instead of absolute values, as we find that the LLM is more reliable at predicting relative rankings than absolute values.
@@
To evaluate MoleVers, we introduce a new benchmark, Molecular Property Prediction in the Wild (MPPW). This benchmark consists of 22 small datasets curated from the ChemBL database (Zdrazil et al., 2024). These datasets, most of which containing 50 or fewer training labels, span a wide range of molecular properties from physical characteristics to biological activities. We standardized the pretraining datasets and data splits to ensure fair comparisons between MoleVers and several state-of-the-art pretrained models. Experimental results show that MoleVers outperforms all baselines in 18 out of the 22 assays and ranks second in the remaining four, while no baseline method consistently ranks in the top two. Moreover, MoleVers achieves state-of-the-art performance on large datasets in the MoleculeNet benchmark (Wu et al., 2018), highlighting the effectiveness of our two-stage pretraining strategy.
@@
In summary, our contributions are: (1) a two-stage pretraining framework with a novel extreme denoising task enabled by our branching encoder and dynamic noise scale sampling in the first stage, and auxiliary predictions based on DFT or LLM in the second stage, (2) a novel branching encoder architecture that facilitates the extreme denoising pretraining, (3) a novel approach to utilize LLM as a molecular property ranker for auxiliary labels generation, and (4) the MPPW benchmark, designed to reflect real-world data limitations. Source code is available at https://github.com/ktirta/MoleVers.
2 Related Work
@@
Deep learning-based molecular property prediction has demonstrated remarkable successes. Early approaches use graph neural networks (GNNs) to learn molecular representations directly from molecular structures (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018). GNNs typically learn molecular representations by updating the node (atom) and edge (bond) features through a series of message passing across neighboring atoms. Recently, popular property prediction benchmarks such as MoleculeNet (Wu et al., 2018) are dominated by transformer-based models (Luo et al., 2022; Zhou et al., 2023; Yang et al., 2024) that leverage self-attention mechanisms to learn long-range interactions between atoms in a molecule.
@@
Parallel to advances in architecture, pretraining has emerged as an effective strategy to improve property prediction performance when labeled data is limited. By pretraining on a large, unlabeled dataset, a model can learn robust and transferable molecular representations that generalize well to a variety of downstream tasks. Various pretraining strategies have been proposed, including masked predictions (Wang et al., 2019; Xia et al., 2023; Zhou et al., 2023; Yang et al., 2024) and contrastive learning (Liu et al., 2022; Xia et al., 2023; Wang et al., 2022). Additionally, denoising atom coordinates and pairwise distance between them (Zaidi et al., 2023; Zhou et al., 2023; Liu et al., 2023) have been shown to lead to strong downstream performance. Pretraining via denoising is equivalent to learning an approximate molecular force field (Zaidi et al., 2023; Liu et al., 2023), which could explain its effectiveness for improving downstream property prediction performance.
Our work is also related to the few-shot molecular property prediction. Previous studies in this area (Ju et al., 2023; Guo et al., 2021; Wang et al., 2021) often formulate the few-shot prediction as an N-way K-shot classification problem, where N classes of molecules are sampled from a dataset, each with K examples. As this formulation is not directly applicable to regression tasks, we focus our discussion in the following sections to studies that follow the pretraining-finetuning paradigm.
@@ 3 Two-Stage Pretraining
Our primary objective is to develop an accurate molecular property prediction model that works in extremely small data regimes. To address this challenge, we propose a two-stage pretraining framework specifically designed to improve the generalization capability of our model, MoleVers. This approach enables accurate property prediction with only a few downstream labels. In the following subsections, we discuss the details of MoleVers including the proposed extreme denoising pretraining, branching encoder architecture, and auxiliary label generation with density functional theory and large language models.
3.1 Stage 1: Masked Atom Prediction and Extreme Denoising
In the first stage of our pretraining framework, we propose to train MoleVers on a large, unlabeled dataset using a combination of masked atom prediction (MAP) (Zhou et al., 2023) and a novel extreme denoising strategy, enabled by our branching encoder and dynamic noise scale sampling. The following subsections provide a detailed discussion of these components.
@@ 3.1.1 Masked Atom Prediction
Inspired by masked token prediction in natural language processing (NLP) (Devlin et al., 2019; Liu et al., 2019; Lewis et al., 2020), masked atom prediction (MAP) involves training a model to predict the correct atom types in a partially-masked molecule. This encourages the model to learn contextual relationship between atom types, capturing how they co-exist in various molecules. Multiple works (Zhou et al., 2023; Xia et al., 2023; Yang et al., 2024) have demonstrated the effectiveness of MAP as a pretraining task, which ultimately leads to better prediction models for the downstream datasets. Following these studies, we use MAP as one of the two pretraining tasks in the first pretraining stage.
3.1.2 Extreme Denoising
@@
Here, we propose extreme denoising, a novel pretraining task that leverages relatively large noise scales to improve model generalization. Our hypothesis is that using higher noise scale values (e.g., $\sigma=10$ instead of $\sigma=1$) exposes the model to a broader set of non-equilibrium molecular configurations, ultimately improving its ability to generalize. However, prior studies (Zhou et al., 2023; Yang et al., 2024; Ni et al., 2024) have reported that larger $\sigma$ values often deteriorate the downstream performance. To address this challenge, we introduce two novel components: a branching encoder architecture which decouple the MAP and denoising pipelines, and the dynamic noise scale sampling.
@@
Decoupling MAP from Denoising. We first motivate the design choice of our branching encoder by examining the complexity differences of the MAP and denoising tasks. In MAP, the model learns to map masked atoms (${\bm{A}}^{\text{mask}}$) to their corresponding atom logits ($\hat{\bm{A}}$), $f({\bm{A}}^{\text{mask}})=\hat{\bm{A}}$, while in coordinate denoising, it learns to map noisy coordinates to their original values, $g(\tilde{{\bm{P}}})=\hat{\bm{P}}$. The MAP function is relatively simpler because it maps a finite set of inputs (atom types) to a relatively compact set of outputs (softmax-normalized logits). In contrast, denoising deals with continuous input and output coordinates, making it more complex as the number of possible mappings is much larger. When a single model handles both MAP and denoising, the overall complexity is dominated by the more challenging denoising task. The downstream performance could then be negatively affected if the model struggles to accurately fit the complex denoising function.
@@
To mitigate this issue, we introduce a branching encoder architecture, shown in Figure 1, that decouples the MAP and denoising pipelines. The branching design ensures that the complexity of the MAP task is only minimally affected by the denoising task. Furthermore, we propose to connect the two encoders with an aggregator module so that information can flow between the two pipelines.
@@
Figure 2: In pretraining stage 2 and finetuning, we keep only the primary encoder to encode the atom and pair distance representations. A prediction head is appended to the model to predict the numerical properties of the molecules, denoted as yreg. LLM-generated ranking labels, denoted as yrank, can be used in the second pretraining stage or during finetuning.
@@
Branching Encoder. Inspired by prior work in NLP which have found masked prediction to often be the most effective pretraining tasks (Lewis et al., 2020; Raffel et al., 2020), we set the MAP encoder as the primary encoder of the model. The primary encoder, shown in Figure 1, will be further pretrained in the second pretraining stage (Figure 2) before being used for downstream predictions.
@@
The branching encoder takes as input the types ${\bm{A}}\in{\mathbb{Z}}^{N}$ and coordinates ${\bm{P}}\in{\mathbb{R}}^{N\times 3}$ of the $N$ atoms in a molecule. Following Zhou et al. (2023) and Yang et al. (2024), each atom type is encoded into atom representation ${\bm{X}}\in{\mathbb{R}}^{N\times C}$ where $C$ is the number of features. The coordinates are transformed into pair distance representation ${\bm{D}}\in{\mathbb{R}}^{N\times N}$. During the first pretraining stage, the atom representations are masked with a ratio of r. We denote the masked atom representations as ${\bm{X}}^{m}$.
@@
To extract the molecule representation ${\bm{F}}$, we feed ${\bm{X}}^{m}$ and ${\bm{D}}$ into the primary MAP encoder $\phi^{p}$. The logits $\hat{\bm{A}}$ that represent the pristine atom types are then predicted with the MAP head $\psi^{p}$,
@@
$\tilde{{\bm{X}}}={\bm{X}}+\bm{\epsilon}_{1},\hskip 10.00002pt\tilde{{\bm{P}}}={\bm{P}}+\bm{\epsilon}_{2},$ (2)
$\bm{\epsilon}_{1}\sim\mathcal{N}(\textbf{0},\sigma^{2}{\bm{I}}_{N\times 1}),\hskip 5.0pt\bm{\epsilon}_{2}\sim\mathcal{N}(\textbf{0},\sigma^{2}{\bm{I}}_{N\times 3}),$ (3)
@@
$\tilde{{\bm{X}}}^{\text{aug}}=[\tilde{{\bm{X}}},\phi^{a}({\bm{F}})],\hskip 10.00002pt{\bm{G}}=\phi^{d}(\tilde{{\bm{X}}}^{\text{aug}},\tilde{{\bm{D}}},\sigma),$ (4)
$(\hat{\bm{X}},\hat{\bm{P}},\hat{\bm{D}})=\psi^{d}({\bm{G}}),$ (5)
@@
Dynamic Noise Scale Sampling. Conventional coordinate-based denoising typically employs a fixed noise scale $\sigma$ throughout the pretraining (Zaidi et al., 2023; Zhou et al., 2023). However, we find that dynamically changing the noise scale during pretraining improves downstream performance. Instead of using a static $\sigma$, we sample it from a uniform distribution, $\sigma\sim\mathcal{U}(0,a)$ where $a>1$. This approach effectively alters the variance of the Gaussian from which the noise is sampled, introducing a bias toward lower noise levels compared to a $\mathcal{N}(0,a)$ distribution (Figure 8). As a result, MoleVers is exposed more frequently to easier, low-noise samples while still ocassionaly encountering challenging, high-noise samples, leading to a more guided learning process. In contrast, a static and low-valued $\sigma$ limits training to only easy samples, whereas a static and high-valued $\sigma$ makes difficult samples more frequent, potentially hindering the model’s convergence.